OpenStudy (anonymous):

You are buying gifts for 10 people. You decide to buy each person either a CD or a DVD. A D costs $12 and a DVD costs $20. A. Let C= the number of CDs you decide to buy. What is an expression in terms of C for the number of DVDs you buy? B. What is an expression in terms of C for the costs of CDs? For the cost of DVDs? C. Write and simplify an expression in terms of C for the total cost of all the gifts you buy. What properties of real numbers did you use to simplify the expression?

OpenStudy (anonymous):

i really need help with this question!!!!

OpenStudy (anonymous):

dvd = D (cost 20) cd = C (cost 12) A) C + D = 10 : C = 10 - D ?? not sure because this expression will tell you how many cd's you bought. D = 10 - C tells you how many dvd's you bought B) cost of CD's -- 12C + 20D = total cost (T) 12C = T - 20D C = (T - 20D)/12 cost of DVD's -- 12C + 20D = total cost 20D = T - 12C D = (T - 12C)/20 C) not sure on this one you might better get a second opinion on this....